OpenStudy (anonymous):

how would I solve 8x-12=4x+24

OpenStudy (phi):

first, add -4x to both sides. (write -4x on both sides) then simplify. what do you get ?

OpenStudy (anonymous):

wait 12-12 and 24-12 then I minus 8x -4x 12x=12

OpenStudy (anonymous):

First subtract both sides by 4x: 4x - 12 = 24 Then add 12 to both sides: 4x = 36 Divide by 4: x = 8

OpenStudy (phi):

can you do it in short steps? something went wrong. 8x-12=4x+24 add -4x

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

8x-12=4x+24 X on one side so minus 4x from both sides 4x-12=24 Then add 12 to both sides 4x=36 Divide both side by 4 x=9

OpenStudy (phi):

8x-12=4x+24 write -4x on both sides, like this 8x-4x -12=4x-4x+24 simplify